OpenStudy (anonymous):

solve for x: log(10-3x)-2logx=0

OpenStudy (anonymous):

log(10-3x) = 2logx log(10-3x) = logx^2 10 - 3x = x^2 x^2 +3x -10 = 0 (x + 5)(x - 2) = 0 x = -5, 2 Eliminate the -5 because you can't take the logarithm of a negative number, and x = 2.

OpenStudy (anonymous):

wow thx for explainin it i get it im doin review assigment so kinda forgot this stuff but its comin back thx